Project

General

Profile

07/30/2014 Minutes

Attendance: Gavin, Jose, Jim, Lisa, Kyle

Jim - Number of noisy FEBs is around 50. Installation process is complete, go back to fix these.
FEBs drop out over the course of a run. Assumed to be noisy. Suggestion that's not the case.
An independent list of FEBs in that case.
Make a list of FEBs present at a start of a run and compare to end of run.

Partition 1 and partition 2 both show diblock 1 and 5 over the last 24 hours.

Point for Joao: All data flagged as bad for the last week - failing the livetime cut.

Near Detector
Plots not updated for 24 hours due to no beam.
Separate set would be useful for cosmic stream.
Jim and Hayes working on getting Ana plots up for Near Detector.

Slice/Track ratio issue. Resolved?

Acceptance issues.
Martin is interested in seeing full DDT Trigger rate plots. Kyle is working on this.